OpenStudy (anonymous):

dividing by a fraction will increase the answer but multiplying by a fractionwill decrease the answer. true or false ?

myininaya (myininaya):

false counterexample: \[3/2 \div 4/3= 3/2 \times 3/4=9/8 but 3/2>9/8 this decreased the answer\]

OpenStudy (anonymous):

depends on the fraction. This is only true if\[\left| fraction \right| <1\] If not, this is false